OpenStudy (anonymous):

What property is shown in the equation? 7ab • 1 = 7ab A. Associative Property of Multiplication B. Commutative Property of Multiplication C. Identity Property of Multiplication D. Reciprocal Property of Multiplication

OpenStudy (dani_rose):

It looks like C.

OpenStudy (anonymous):

can u help with some more

OpenStudy (dani_rose):

Sure

rebeccaxhawaii (rebeccaxhawaii):

yes C

OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. What property is shown by the equation? Column A Column B 1. –24 • 0 = 0 2. 1/16 • (16)=1 3. –18 • (–1) = 18 4. 5 ÷ (–1) = –5 A. Reciprocal Property of Multiplication B. Identity Property of Division C. Negative One Property of Multiplication D. Zero Property of Multiplication E. Negative One Property of Division

OpenStudy (kittiwitti1):

first one: -24 • 0 what does it look like? hint: there's a zero.

OpenStudy (anonymous):

c?

OpenStudy (kittiwitti1):

I said "there's a zero". does C talk about a zero at all? lol

OpenStudy (anonymous):

i ment d

OpenStudy (kittiwitti1):

oh. okay then, good job :D

OpenStudy (kittiwitti1):

ok, second one. 1/16 • 16 = 1 do you know what a reciprocal is?

OpenStudy (anonymous):

kinda my teacher didn't explane what it was very well

OpenStudy (kittiwitti1):

ok well a reciprocal is basically 1/(fraction) which flips the fraction:\[\frac{a}{b}\text{ reciprocal }\rightarrow\frac{1}{\frac{a}{b}}=1\times\frac{b}{a}=\frac{b}{a}\]

OpenStudy (kittiwitti1):

so what's the reciprocal of 1/16 ?? hint: a = 1, b = 16...

OpenStudy (anonymous):

16

OpenStudy (kittiwitti1):

yes :-) so what is the answer for #2?

OpenStudy (kittiwitti1):

everything ok?

OpenStudy (anonymous):

yah

OpenStudy (kittiwitti1):

alright what do you think the answer is for #2 :-)

OpenStudy (anonymous):

a

OpenStudy (kittiwitti1):

yes! good job, that's correct ^-^ ok, on to the next one

OpenStudy (anonymous):

yaaa

OpenStudy (kittiwitti1):

#3: -18 • -1 = 18 hint: you're multiplying -18 by -1 ... lol

OpenStudy (anonymous):

c

OpenStudy (kittiwitti1):

oh random fyi, negative times negative cancels out and gives you positive solution :-)

OpenStudy (anonymous):

yup i tack screen shots of info i need got a chart for that stuff

OpenStudy (anonymous):

and am i right

OpenStudy (anonymous):

hello

OpenStudy (kittiwitti1):

sorry I gtg, mom screaming at me and crying (idk???) o_e

OpenStudy (kittiwitti1):

had an argument =_=

OpenStudy (anonymous):

is everything ok

OpenStudy (kittiwitti1):

yeah we're ok, don't worry lol hope you got the rest of the answers! :-)